OpenStudy (1645323):

how do you solve 4^(x+5)= (1/8)^(2x-8)

OpenStudy (aum):

Take log to the base 2 on both sides.

OpenStudy (1645323):

what would the result be? My teacher says to put 1/8 to 4^-1.5 and solve but I don't know how

OpenStudy (anonymous):

Is this your equation?\[4^{(x+5)}=\frac{ 1 }{ 8 }^{(2x-8)}\]

OpenStudy (aum):

You can solve it without logs also. The base on the left hand side is 4. The base on the right hand side is 1/8. Make both sides have base 2. 4 = 2^2 1/8 = 1/(2^3) = 2^(-3) \[ \large 4^{x+5}= \left(\frac 18 \right)^{2x-8} \\ \large (2^2)^{x+5} = (2^{-3})^{2x-8} \\ \large 2^{2x+10} = 2^{-6x+24} \\ 2x + 10 = -6x + 24 \\ 8x = 14 \\ x = 14/8 = 7/4 \]

OpenStudy (anonymous):

Yup!\[x = \frac{ 7 }{ 4 }\]

OpenStudy (1645323):

Ok I didn't realize that you went down to 2 on the fraction and i ddin't know you could mulitply the power of (2^2) but thanks!

OpenStudy (aum):

You are welcome.